100 cm3 of solution containing 10 g of pure H2S04 were mixed with an equal volume of a solution containing 15 g of HCL. The mixture was stirred to make it uniform. If there was no volume change during mixing, how many cubic cm of 1 M sodium hydroxide would be needed to neutralize 25 cm3 of the mixture
1
Expert's answer
2020-12-09T06:37:54-0500

V=100 + 100 = 200 ml.

n1= 10 x /98 = 0,102040816 moles. H2SO4

m2= 15/36.5 = 0,410958904 moles. HCl.

n(total)= 2 x 0,102040816 + 0,410958904 = 0,615040536 moles of NaOH.

Cm=n/V.

V= n/Cm.

25 x 0.615040536/200 = 0,076880067 moles of NaOH.

V= 0,076880067/1 = 0,076880067 L.

76.880067 ml.